Brigham and Houston just paid a dividend of $2.00 a share. The dividend is expected to grow at 25% a year for the next 3 years and then at a rate of 6% a year thereafter. What is the expected dividend per share for year 3? Group of answer choices
Answer #1

Expected dividend per share for year 3=Last Dividend*(1+growth rate)^3=2*1.25^3=3.91